Integrate w.r.t to x using substitution
(A) ∫ x^3√x^4+1 dx from 1 to 0
(B) tan^7x sec^2x dx from pi by 4 to 0
(C) sec^2 (cosx) sinx dx
1
Expert's answer
2020-03-10T10:06:21-0400

"A.) \int_1^0 x^3 \sqrt{x^4 + 1} \, dx"

Let us assume that:

"x^4 + 1 = u \\
\Rightarrow 4x^3 \, dx = du"

"\Rightarrow x^3 \, dx =\frac{du}{4}"

When:

"x = 1"

Then:

"u = 1^4 + 1 = 2"

Again, when:

"x = 0"

Then:

"u = 0^4 + 1 = 1"

Substituting "u = x^4 + 1" into the integration and integrating with respect to "du" , we have:

"= \frac{1}{4} \int_2^1 \sqrt{u} \, du"

"= \frac{1}{4} \int_2^1 u^{\frac{1}{2}} \, du"

"= \frac{1}{4} \left[ \frac{u^{\frac{3}{2}}}{\frac{3}{2}} \right]_2^1" "\hspace{1 cm} \left[ \because \int x^n \, dx = \frac{x^{n+1}}{n+1} \right]"

"= \frac{1}{6} \left[ u^{\frac{3}{2}} \right]_2^1"

"= \frac{1}{6} \left[ 1^{\frac{3}{2}} - 2^{\frac{3}{2}} \right] \\
= \frac{1 - 2\sqrt{2}}{6}"

"\\"

B.) "\int_{\frac{\pi}{4}}^0 \tan^7 (x) \sec^2 (x) \, dx"

Let us assume that:

"tan (x) = u \\
\Rightarrow sec^2 (x) \, dx = du"

When:

"x = \frac{\pi}{4}"

Then:

"u = \tan (\frac{\pi}{4}) = 1"

Again, when:

"x = 0"

Then:

"u = \tan (0) = 0"

Substituting "u = \tan (x)" into the integration and integrating with respect to "du" , we have:

"= \int_1^0 u^7 \, du \\
= \frac{1}{8} \left[ u^8 \right]_1^0" "\hspace{1 cm} \left[ \because \int x^n \, dx = \frac{x^{n+1}}{n+1} \right]"

"= - \frac{1}{8}"

"\\"

C.) "\int \sec^2 (\cos(x)) \sin (x) \, dx"

Let us assume that:

"\cos(x) = u \\
\Rightarrow - \sin (x) \, dx = du"

Substituting "u = \cos (x)" into the integration and integrating with respect to "du" , we have:

"= - \int sec^2 (u) \, du \\
= - tan (u) + C \hspace{1 cm}" "\left[ \because \int \sec^2 (x) \, dx = \tan (x) + C \right]"

Where C is a constant of integration.

Undo substitution and we have:

"= - \tan (\cos(x)) + C" "\hspace{1 cm} \left[ \because u = \cos(x) \right]"